OpenStudy (anonymous):

how do you get log cot 0.1130?

OpenStudy (cggurumanjunath):

first find cot 0.1130 in calculator and then find log of the obtained result.

OpenStudy (jack1):

just to clarify though: cot = cotangent = 1/tan so if you want to find cot 45 = 1/tan 30 = 1/ 0.5774 = 1.732 ish ( exact ans = sqrt 3 as tan 30 = 1/ sqrt 3)

OpenStudy (jack1):

*mistake i meant cot 30 (not cot 45)